"Looking Closely at Cost and Competition"

Watch this video, Revenue, Profits, and Price: Crash Course Economics #24, to help you prepare for this week’s  discussion options.  Only ONE OPTION is required!

Option #1: Reply to the following prompts by using the company for which you currently work, a business with which you’re familiar, or the dream business you want to start:

  • Do you think it’s easy for your selected business to enter this same industry?
  • What are some key fixed, variable, implicit, and/or opportunity costs?

Discuss with your peers:

  • Read one of your peer’s posts from any discussion option and share another idea for a type of cost, profit, etc.

Option #2: Consider the table below.

  • What is the accounting cost and accounting profit for this business?
  • What is the economic cost and economic profit?
  • Why do these estimates of costs and profit differ?

Discuss with your peers:

  • Read one of your peer’s posts from any discussion option and share another idea for a type of cost, profit, etc.
Total Revenue $600,000.00
Itemized Costs:
Total annual labor cost (the business owner does not draw a salary)   $160,000.00
Total annual cost of materials used in production $250,000.00
Annual expenditure on rent, utilities, taxes, insurance, and misc. expenses $90,000.00
The business owners' previous annual salary when he worked for someone else $85,000.00
The business owners' annual interest income on past savings before using it up to start this business $15,000.00

Option #3: Suppose an auto maker is able to sell one line of cars for $30,000 each. All cars are identical and the price can not be raised due to competition. Now consider the output levels and total costs in the schedule below and answer the following questions:

  • Does it make sense for this auto maker to produce 11 cars and why?
  • Does it make sense for this auto maker to produce 12 cars and why?

Discuss with your peers:

  • Read one of your peer’s posts from any discussion option and share another idea for a type of cost, profit, etc.
Output level Total Cost Average Cost per car
10 cars $200,000.00 $20,000.00
11 cars $227,500.00 $20,681.82
12 cars $260,000.00 $21,666.67
Answer #1

Answer

Option 2:

1)Accounting cost =total labor cost +total material cost +Annual cost =160000+250000+90000=$500000

Accounting Profit = Revenue - Accounting cost =600000-500000=$100000

2)

Opportunity Cost =salary + interest income =85000+15000=$100000

Economic cost = Accounting Cost + Opportunity cost =500000+100000=$600000

Economic Profit =revenue -Economic cost=600000-600000=$0

Since economic profit is Zero so he is earning normal profit.

3)

Since in economic cost and profit calculation we are also including opportunity cost.so cost increase and profit decreases.